The Brisbane Manufacturing Company produces a single model of a CD player. Each player is sold for $195 with a resulting contribution margin of $76. Brisbane's management is considering a change in its quality control system. Currently, Brisbane spends $42,000 a year to inspect the CD players. An average of 2,200 units turn out to be defective - 1,760 of them are detected in the inspection process and are repaired for $75. If a defective CD player is not identified in the inspection process, the customer who receives it is given a full refund of the purchase price. The proposed quality control system involves the purchase of an x-ray machine for $200,000. The machine would last for four years and would have salvage value at that time of $21,000. Brisbane would also spend $460,000 immediately to train workers to better detect and repair defective units. Annual inspection costs would increase by $21,000. This new control system would reduce the number of defective units to 390 per year. 335 of these defective units would be detected and repaired at a cost of $41 per unit. Customers who still received defective players would be given a refund equal to 120% of the purchase price. Questions 1 & 2 1. What is the Year 2 cash flow if Brisbane keeps using its current system? 2. What is the Year 2 cash flow if Brisbane replaces its current system? Questions 3 & 4 3. Assuming a discount rate of 6%, what is the net present value if Brisbane keeps using its current system? 4. Assuming a discount rate of 6%, what is the net present value if Brisbane replaces its current system?
Answer #1

cash flow for current system

Total defective units-detected

=2200-1760

=440units

year 2 cash flow

Annual inspection cost -repairing cost *units repaired- refunded units

= -42000 -(75*1760)-440*195

=-42000-132000-85800

= -259800$

new system cash flow in year 2

annual inspection cost -repairing cost*units repaired - refunded units

= -[42000+21000increase]-335*41-[390-335]*195*120%

= -63000-13735-12870

= -89605

NPV

CURRENT SYSTEM

YEAR cash flow pv factor at 6% presnt value
1-4 -259800 3.46511 -900236

new system

YEAR cash flow pv factor at 6% presnt value
1-4 -89605 3.46511 -310491[89605*3.46511]
0 -200000 1 -200000
0 -460000 1 -460000
4 21000 0.7921 16634
NET PRESENT VALUE .-953857